# FIARSE: Model-Heterogeneous Federated Learning via Importance-Aware Submodel Extraction

### v1

@@ C.2 Preliminary
$\quad+4K\left(1+36\eta_{l}^{2}L^{2}(1+K)\right)\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}$ (62)
@@
C.3 Main Proof
Theorem C.5.
Suppose that Assumption 4.1, 4.2 and 4.3 hold. We define $F(\tilde{\boldsymbol{x}})\overset{\triangle}{=}F\left(\tilde{\boldsymbol{x}},\boldsymbol{1}^{N\times d}\right)$, and $F(\tilde{\boldsymbol{x}})\geq F_{*}$ for all $\tilde{\boldsymbol{x}}\in\mathbb{R}^{d}$. Let the local learning rate satisfy $\eta_{l}\leq\min\left(\frac{1}{2L\sqrt{K(K+1)}},\frac{1}{6L\sqrt{(K+1)A}},\frac{\eta_{s}}{9L},\frac{1}{12L\sqrt{KN}},\frac{A}{32KNL\eta_{s}}\right)$. Denote $T$ as the total communication rounds. Therefore, the convergence rate of FIARSE for non-convex objectives should be
Proof.
@@ C.3 Main Proof
Apparently, the parameters are updated only when the submodels hold the counterpart. In this means, the definition for $\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)$ is
@@
$\quad\mathbb{E}\left\langle\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right),\tilde{\boldsymbol{x}}_{t+1}-\tilde{\boldsymbol{x}}_{t}\right\rangle=\sum_{j=0}^{n}\mathbb{E}\left\langle\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right),\tilde{\boldsymbol{x}}_{t+1}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\right\rangle$ (68)
$=\sum_{j=0}^{n}\mathbb{E}\left\langle\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right),-\eta_{s}\eta_{l}\cdot\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\rangle$ (69)
$=-\eta_{s}\eta_{l}K\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\mathbb{E}\left\langle\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right),\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\rangle$ (70)
@@
$\quad+\frac{\eta_{s}\eta_{l}K}{2}\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (72)
where the last equation is built upon $\left\langle\boldsymbol{a},\boldsymbol{b}\right\rangle=-\frac{1}{2}\left\|\boldsymbol{a}\right\|_{2}^{2}-\frac{1}{2}\left\|\boldsymbol{b}\right\|_{2}^{2}+\frac{1}{2}\left\|\boldsymbol{a}-\boldsymbol{b}\right\|_{2}^{2}\leq-\frac{1}{2}\left\|\boldsymbol{a}\right\|_{2}^{2}+\frac{1}{2}\left\|\boldsymbol{a}-\boldsymbol{b}\right\|_{2}^{2}$. In view that
@@
we simplify Equation (72) with the conclusion from Lemma C.2:
@@
$\leq$ $-\frac{\eta_{s}\eta_{l}KA}{2N}\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}+\frac{\eta_{s}\eta_{l}K}{2}\left(4A\delta^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+72KA\eta_{l}^{2}L^{2}\left(\sigma^{2}+\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\right)\right)$ (75)
@@
Based on Theorem C.5, when we apply local and global learning rates as described in Theorem 4.5, we can obtain the desired conclusion that the proposed FIARSE can converge to a stationary point at a rate of $O(1/\sqrt{T})$.
@@
Table 4: Test accuracy under four different submodel sizes on different datasets, and 20% out of 100 clients participate in the training at each round. To be more specific, the columns from “Local” to “Model (1.0)” evaluate the test accuracy on the local test datasets, while “Global” evaluates the test accuracy of the global model on the global test dataset.
  Part. Ratio Method CIFAR-10 CIFAR-100

### v2

@@ C.2 Preliminary
$\bar{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}$ $=\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)-\nabla F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right),$ (71)
$\ddot{\boldsymbol{g}}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}$ $=\nabla F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right).$ (72)
Since
$\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}=\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}+\bar{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}+\ddot{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}+\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right),$ (73)
where $\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}$ is defined in Equation (50), we bound $\left\|\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ by splitting it into four terms:
$\left\|\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (74)
$\leq$ $4\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}+4\left\|\bar{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}\right\|_{2}^{2}+4\left\|\ddot{\boldsymbol{g}}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}+4\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}.$ (75)
Therefore, to bound Equation (70), we should analyze the following inequality, i.e.,
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (76)
$\leq$ $4\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}$ (77)
$+4K\cdot\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\bar{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}\right\|_{2}^{2}$ (78)
$+4K\cdot\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\ddot{\boldsymbol{g}}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (79)
$+4K\cdot\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}.$ (80)
There are four terms in the above inequality, i.e., Equation (77), Equation (78), Equation (79), and Equation (80). Subsequently, we analyze these four terms one by one.
$\bullet$ For Equation (77), we apply Lemma C.2 and obtain that
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}\leq\frac{A}{N}\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}.$ (81)
In Lemma C.4, we mention that
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}=\sum_{i\in[N]}\sum_{k=0}^{K-1}\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[0:\gamma_{i}](i)}_{t,k}\right\|_{2}^{2};\quad\left\|\tilde{\boldsymbol{g}}^{[0:\gamma_{i}](i)}_{t,k}\right\|_{2}^{2}\leq L^{2}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}.$ (82)
Therefore, Equation (77) is further bounded by
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}\leq\frac{AL^{2}}{N}\sum_{i\in[N]}\sum_{k=0}^{K-1}\mathbb{E}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}.$ (83)
$\bullet$ Similarly, Equation (78) is bounded for
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\bar{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}\right\|_{2}^{2}\leq\frac{A}{N}\sum_{i\in[N]}\mathbb{E}\left\|\bar{\boldsymbol{g}}^{[0:\gamma_{i}](i)}_{t}\right\|_{2}^{2}.$ (84)
In view that the non-zero part of $\nabla F_{i}^{[0:\gamma_{i}]}(\tilde{\boldsymbol{x}}_{t})$ is equivalent to $\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)$, we apply Assumption 5.3 and attain that
$\left\|\bar{\boldsymbol{g}}^{[0:\gamma_{i}](i)}_{t}\right\|_{2}^{2}\leq\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}.$ (85)
Therefore, the bound of Equation (78) is
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\bar{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}\right\|_{2}^{2}\leq A\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}.$ (86)
$\bullet$ For Equation (79), we apply $\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\leq 1$ and Assumption 5.2 and obtain
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\ddot{\boldsymbol{g}}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}\leq\sigma^{2}$ (87)
$\bullet$ With $\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\leq 1$, Equation (80) is bounded and simplified for
$\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\leq\sum_{j=0}^{n}\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}=\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}$ (88)
In conclusion, Equation (70) is bounded by
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (89)
$\leq$ $\frac{4AL^{2}}{N}\sum_{i\in[N]}\sum_{k=0}^{K-1}\mathbb{E}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+4KA\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+4K\sigma^{2}+4K\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}$ (90)
By applying Lemma C.3, we further simplify the inequality for
$\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (91)
$\leq$ $4KA\left(1+36\eta_{l}^{2}KL\right)\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+4K\left(1+36\eta_{l}^{2}KAL\right)\sigma^{2}+4K\left(1+36\eta_{l}^{2}KAL\right)\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}$ (92)
@@
C.3 Main Proof of Theorem 5.4
@@
Similar to Lemma C.5, we define $\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}$ to be
$\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}=\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right).$ (96)
Therefore,
$\textsf{Agg}_{i\in\mathcal{A}}\left(\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right)=\bigcup_{j\in[n]}\frac{1}{\left|\mathcal{A}_{\gamma_{j}^{\prime}}\right|}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}.$ (97)
In the proposed algorithm, the parameters are updated only when the submodels hold the counterpart. In this means, the definition for $\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)$ is
@@
$\quad\mathbb{E}\left\langle\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right),\tilde{\boldsymbol{x}}_{t+1}-\tilde{\boldsymbol{x}}_{t}\right\rangle=\sum_{j=0}^{n}\mathbb{E}\left\langle\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right),\tilde{\boldsymbol{x}}_{t+1}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}-\tilde{\boldsymbol{x}}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\right\rangle$ (99)
$=\sum_{j=0}^{n}\mathbb{E}\left\langle\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right),-\eta_{s}\eta_{l}\cdot\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\rangle$ (100)
$=-\eta_{s}\eta_{l}K\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\mathbb{E}\left\langle\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right),\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\rangle$ (101)
@@
$\quad+\frac{\eta_{s}\eta_{l}K}{2}\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (103)
where the last equation is built upon $\left\langle\boldsymbol{a},\boldsymbol{b}\right\rangle=-\frac{1}{2}\left\|\boldsymbol{a}\right\|_{2}^{2}-\frac{1}{2}\left\|\boldsymbol{b}\right\|_{2}^{2}+\frac{1}{2}\left\|\boldsymbol{a}-\boldsymbol{b}\right\|_{2}^{2}\leq-\frac{1}{2}\left\|\boldsymbol{a}\right\|_{2}^{2}+\frac{1}{2}\left\|\boldsymbol{a}-\boldsymbol{b}\right\|_{2}^{2}$. Lemma C.2 mentions that
@@
and we simplify Equation (103) with the conclusion from Lemma C.4:
@@
$\leq$ $-\frac{\eta_{s}\eta_{l}KA}{2N}\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}+\frac{\eta_{s}\eta_{l}K}{2}\left(4A\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+72KA\eta_{l}^{2}L^{2}\left(\sigma^{2}+\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\right)\right)$ (106)
@@
Based on Theorem 5.4, when we apply local and global learning rates as described in Corollary 5.5, we can obtain the desired conclusion that the proposed FIARSE can converge to a stationary point at a rate of $O(1/\sqrt{T})$.
@@
Table 4: Test accuracy under four different submodel sizes on different datasets, and 20 out of 100 clients participate in the training at each round. To be more specific, the columns from “Local” to “Model (1.0)” evaluate the test accuracy on the local test datasets, while “Global” evaluates the test accuracy of the global model on the global test dataset.
  Method CIFAR-10 CIFAR-100